OpenStudy (anonymous):

Solve the equation. x^2 = 150

DivineSolar (divinesolar):

Your last question is wrong by the way.

DivineSolar (divinesolar):

The other guy gave you an incorrect answer.

OpenStudy (anonymous):

Really? what the

OpenStudy (amorfide):

divine before telling others are wrong learn to read, 121/4 is 30.25 I told him to square root 30.25

DivineSolar (divinesolar):

Wrong.

OpenStudy (anonymous):

4x2 = 121 (Points : 1) 30.25, −30.25 5.5, −5.5 11, −11 2.75, −2.75

DivineSolar (divinesolar):

+underline sqrt(121/4), x = 5.5, -5.5

DivineSolar (divinesolar):

See you dont just divide it.

OpenStudy (amorfide):

you are saying the exact same thing as i said...

DivineSolar (divinesolar):

See but you have solved it wrong you just divided it.

OpenStudy (amorfide):

divide 121 by 4 then square root your answer...

OpenStudy (anonymous):

can you explain it instead of giving answer

DivineSolar (divinesolar):

Yeah sure.

DivineSolar (divinesolar):

Step 1 : divide both sides by 4 \[\frac{ 4x^2 }{ 4 } = \frac{ 121 }{ 4 } so x^2 = \frac{ 121 }{ 4 }\]

DivineSolar (divinesolar):

Step 2: take the square root \[x = \pm \]

DivineSolar (divinesolar):

sqrt

DivineSolar (divinesolar):

after the plusunderline aka PM

DivineSolar (divinesolar):

in front of the sqrt is \[(\frac{ 121 }{ 4 }) \]

DivineSolar (divinesolar):

which gives you x = 5.5, -5.5

DivineSolar (divinesolar):

See you find the sqrt of 30.25

DivineSolar (divinesolar):

which is 5.5

DivineSolar (divinesolar):

Which gives you 5.5, -5.5 as your final answer.

DivineSolar (divinesolar):

Now do you see how his answer is wrong @Mr1337 ?

OpenStudy (anonymous):

Thanks can you help me out with this question ?

DivineSolar (divinesolar):

Yeah sure just re-post it below my comment here.

DivineSolar (divinesolar):

Never mind its at the top of the page.

DivineSolar (divinesolar):

There is only one step to this which is take the square root.

DivineSolar (divinesolar):

What is 150's square root?

OpenStudy (anonymous):

12

DivineSolar (divinesolar):

12.24

DivineSolar (divinesolar):

Don't remove the decimal unless it asks. so your answer would be 12.247449, - 12.247449

OpenStudy (anonymous):

yeah i got that

DivineSolar (divinesolar):

but your teacher would probably want it more clean so, 12.24, -12.24

OpenStudy (anonymous):

http://gyazo.com/c03fa48920a1036d3c42442844ac200f

OpenStudy (anonymous):

those are the answers

DivineSolar (divinesolar):

What the...

DivineSolar (divinesolar):

Let me see what i can do...

OpenStudy (anonymous):

yeah that's why i was really confused

DivineSolar (divinesolar):

I would go with the first answer. Considering i just solve them here is how i got it -> http://www.tiger-algebra.com/drill/x2=150/

OpenStudy (amorfide):

\[x ^{2}=150\] square root both sides to get \[x=\sqrt{150}\] find a square number that goes into 150, the highest square number in my mind is 25 how many times does 25 go into 150? 6 times you rewrite \[x=\sqrt{150}\] to be \[x=\sqrt{25}\sqrt{6}\] \[\sqrt{25}=\pm5\] which gives \[\pm5\sqrt{6}\]

DivineSolar (divinesolar):

As i already you just confirmed the first answer i suggested :P

OpenStudy (amorfide):

i know just thought I would help him with a simpler layout

DivineSolar (divinesolar):

Yeah that is simpler but more complex and thought out makes it easier. Since it is broken down into layers. But good job anyways.